OpenStudy (anonymous):

A spherical snowball melts so that its radius decreases from 6 inches to 5.8 inches. a) Find the exact change in volume of the snowball. b) Find the approximate change in the volume of the snowball.

OpenStudy (anonymous):

volume of a sphere is \(\frac{2}{3}\pi r^3\) you have to compute \[\frac{4}{3}\pi \times 6^3-\frac{4}{3}\pi \times (5.8)^3\]

OpenStudy (anonymous):

use a calculator since it says approximate

OpenStudy (anonymous):

actually i got that exactly backwards it says approximate change from 6 to 5.8 so it should be the negative of what i wrote above

OpenStudy (anonymous):

how i find the exact change?

OpenStudy (anonymous):

Use a calculator for exact.

OpenStudy (anonymous):

I used one for the approximate too thought